OpenStudy (anonymous):

WALK ME THROUGH THIS PROBLEM AND I WILL MEDAL/FAN! Solve: The quantity 3 x minus 15 divided by 2 = 4x 11 6 −3 −15

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[\frac{3x-15}{2}=4x\]?

OpenStudy (anonymous):

yes ma'am!!

OpenStudy (misty1212):

multiply both sides by 2 to get rid of annoying fraction on the left what do you get?

OpenStudy (anonymous):

3x-15=8x?

OpenStudy (misty1212):

yes

OpenStudy (misty1212):

now to get all the variables on one side, subtract \(3x\) from both sides (that gets rid of the \(3x\) on the left what do you get?

OpenStudy (anonymous):

-15=5x? ^ since we subtracted 3x rather than adding it

OpenStudy (misty1212):

yes

OpenStudy (misty1212):

last step, to get \(x\) by itself, divide both sides by \(5\)

OpenStudy (anonymous):

-3=x?

OpenStudy (misty1212):

yes as always,go with C

OpenStudy (anonymous):

thanks!!!!

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]